OpenStudy (anonymous):

What is the sum of the geometric sequence 3, 15, 75, … if there are 7 terms 39,062 58,593 195,312 292,968

OpenStudy (solomonzelman):

at first, what is the common ratio, and the first term?

OpenStudy (solomonzelman):

The formula for the sum is, \(\Large\color{black}{ \displaystyle {\rm \LARGE S}_{n}=\frac{a_1\left(1-r^{n}\right)}{1-r}}\)

OpenStudy (solomonzelman):

where 'n' is number of terms, \(\Large\color{black}{a_1}\) is the first term, and 'r' is the common ratio.

OpenStudy (solomonzelman):

I know how to deduce it though

OpenStudy (solomonzelman):

algebraically, without this formula

OpenStudy (solomonzelman):

no, you do he formula, and I will post my thingy

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

from 3 we are multiplying times 3, so `(2nd term)+(3rd term)` => will form sum number divisible by 10. `(4th term)+(5th term)` => will form sum number divisible by 10. `(6th term)+(7th term)` => will form sum number divisible by 10. like 15 + 75 = 90 (has 0 at the end) and so will the other 2 sums be, `4th+5th` AND `6th+7th` have 0 at the end. and therefore, the sum of all 6 last terms (not including the first) will get you a number that is divisible by 10. And if so, then after you add 3 to this number, `_ _ _ , _ _0 ` + 3 = `_ _ _ , _ _3 ` THE LAST DIGIT IN THE ANSWER IS 3.

OpenStudy (anonymous):

I got 58,593

OpenStudy (solomonzelman):

good

OpenStudy (solomonzelman):

my first line has a typo

OpenStudy (solomonzelman):

from 3 we are multiplying times 5, so (times 5, not times 3)

OpenStudy (solomonzelman):

I will depart for a small chess game for now.... excuse that